OpenStudy (anonymous):

Find the sum of the series 20+27+34+41+...+664+671

OpenStudy (unklerhaukus):

20+671=691 27+664=691 can you use this pattern some how

OpenStudy (anonymous):

I can't remember how to do this on paper, but I find this would be easier to do with a program.

OpenStudy (callisto):

First term = 20 Common difference = 27-20 = 7 no. of terms = (671 - 20)/7 +1 = ??? Them use the formula Sum = (n/2) (a+l), where a = first term, l= last term and n= no. of terms

OpenStudy (lgbasallote):

i would suggest you find the number of terms then use the formula for sum of arithmetic series \[\large S_n = \frac{n(A_1 + A_n)}{2}\] but unkle's could work

OpenStudy (callisto):

To make it clear no of terms \[=\frac{ (671 - 20)}{7} +1\]